# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ 2 Quantum torus algebra and the representation at level-$k$
3 Representation of a q-deformed Lorentz algebra
@@ 3.1 The representation
Based on the representation $(\mathcal{H},\rho)$ of the quantum torus algebra and by the linear combinations of the operators $\bm{x}_{\alpha\beta}$, we obtain a representation of $U_{{\bf q}}(sl_{2})$ on $\mathcal{H}$. Indeed, we define a family of operators $E_{\lambda},F_{\lambda},K_{\lambda},K_{\lambda}^{-1}$ by
$K_{\lambda}$ $=\bm{x}_{-1,0},\qquad K_{\lambda}^{-1}=\bm{x}_{1,0},\qquad E_{\lambda}=-\frac{1}{{\bf q}-{\bf q}^{-1}}\bm{x}_{-1,1},$ (3.1)
$F_{\lambda}$ $=-\frac{1}{{\bf q}-{\bf q}^{-1}}\left[\left(\lambda+\lambda^{-1}\right)\bm{x}_{1,-1}+\bm{x}_{3,-1}+\bm{x}_{-1,-1}\right],$ (3.2)
@@
where the parameter $\lambda\in\mathbb{C^{\times}}$. It is straight-forward to check that $E_{\lambda},F_{\lambda},K_{\lambda},K_{\lambda}^{-1}$ satisfy the commutation relation of $U_{{\bf q}}(sl_{2})$  [^2] See the convention in [4] up to an inverse of $\mathbf{q}$.:
$K_{\lambda}E_{\lambda}={\bf q}^{-1}E_{\lambda}K_{\lambda}$ $,\qquad K_{\lambda}F_{\lambda}={\bf q}F_{\lambda}K_{\lambda},\qquad\left[E_{\lambda},F_{\lambda}\right]=\frac{K_{\lambda}^{2}-K_{\lambda}^{-2}}{{\bf q}-{\bf q}^{-1}}.$
@@
Similarly, we define the tilded operators
$\tilde{K}_{\lambda}$ $=\tilde{\bm{x}}_{-1,0},\qquad\tilde{K}_{\lambda}^{-1}=\tilde{\bm{x}}_{1,0},\qquad\tilde{E}_{\lambda}=-\frac{1}{\tilde{{\bf q}}-\tilde{{\bf q}}^{-1}}\tilde{\bm{x}}_{-1,1},$
$\tilde{F}_{\lambda}$ $=-\frac{1}{\tilde{{\bf q}}-\tilde{{\bf q}}^{-1}}\left[\left(\tilde{\lambda}+\tilde{\lambda}^{-1}\right)\tilde{\bm{x}}_{1,-1}+\tilde{\bm{x}}_{3,-1}+\tilde{\bm{x}}_{-1,-1}\right],$
They satisfy the commutation relation of $U_{\tilde{{\bf q}}}(sl_{2})$:
$\tilde{K}_{\lambda}\tilde{E}_{\lambda}=\tilde{{\bf q}}^{-1}\tilde{E}_{\lambda}\tilde{K}_{\lambda},\qquad\tilde{K}_{\lambda}\tilde{F}_{\lambda}=\tilde{{\bf q}}\tilde{F}_{\lambda}\tilde{K}_{\lambda}$ $,\qquad\left[\tilde{E}_{\lambda},\tilde{F}_{\lambda}\right]=\frac{\tilde{K}_{\lambda}^{2}-\tilde{K}_{\lambda}^{-2}}{\tilde{{\bf q}}-\tilde{{\bf q}}^{-1}}.$
It shows that $\mathcal{H}$ carries a $\star$-representation of $U_{{\bf q}}(sl_{2})\otimes U_{\tilde{{\bf q}}}(sl_{2})$ labelled by the parameters $(\lambda,\tilde{\lambda})$ with $\lambda,\tilde{\lambda}\in\mathbb{C^{\times}}$. When $\lambda^{*}=\tilde{\lambda}$, the tilded operators and untilded operator are related by the Hermitian conjugate
$\tilde{K}_{\lambda}=K_{\lambda}^{\dagger},\qquad\tilde{E}_{\lambda}=-E_{\lambda}^{\dagger},\qquad\tilde{F}_{\lambda}=-F_{\lambda}^{\dagger},$
as the representation of the $\star$-structure. The minus sign is due to $\bf{q}^{*}=\tilde{\bf{q}}^{-1}$. The representation is irreducible because the representation of the Weyl algebra is irreducible [^3] The generators of the Weyl algebra relate to the $U_{{\bf q}}(sl_{2})$ generators by $\bm{x}=\bm{x}_{1,0}=K_{\lambda}^{-1}$, $\bm{y}=\bm{x}_{0,1}=q^{-1}\bm{x}_{1,0}\bm{x}_{-1,1}=-(\mathbf{q}-\mathbf{q}^{-1})q^{-1}K_{\lambda}^{-1}E_{\lambda}$. . We denote this representation by $\pi_{\lambda,\tilde{\lambda}}$ 444 The above representation is motivated by generalizing the representations studied in e.g. [2, 3, 1] for the modular double of ${U}_{q}(\mathrm{SL}(2,\mathbb{R}))$, although here $(\mathbf{q},\tilde{\mathbf{q}})\to(e^{2\pi ib^{2}},e^{2\pi ib^{-2}})$ when $k=1$ or $(\mathbf{q},\tilde{\mathbf{q}})\to(-e^{\pi ib^{2}}$ and $-e^{\pi ib^{-2}})$ when $k=2$, whereas the modular deformation parameter there is $q=e^{\pi ib^{2}}$ and $\tilde{q}=e^{\pi ib^{-2}}$..
The comultiplication $\Delta$ of $U_{{\bf q}}(sl_{2})\otimes U_{\tilde{{\bf q}}}(sl_{2})$ is defined by
$\Delta K$ $=K\otimes K,\qquad\Delta E=E\otimes K+K^{-1}\otimes E,\qquad\Delta F=F\otimes K+K^{-1}\otimes F,$
$\Delta\tilde{K}$ $=\tilde{K}\otimes\tilde{K},\qquad\Delta\tilde{E}=\tilde{E}\otimes\tilde{K}+\tilde{K}^{-1}\otimes\tilde{E},\qquad\Delta\tilde{F}=\tilde{F}\otimes\tilde{K}+\tilde{K}^{-1}\otimes\tilde{F}.$
@@
$Q_{\lambda}$ $=\left({\bf q}-{\bf q}^{-1}\right)^{2}F_{\lambda}E_{\lambda}-\mathbf{q}^{-1}K_{\lambda}^{2}-{\bf q}K_{\lambda}^{-2}$
$\tilde{Q}_{\lambda}$ $=\left(\tilde{{\bf q}}-{\bf\tilde{q}}^{-1}\right)^{2}\tilde{F}_{\lambda}\tilde{E}_{\lambda}-\tilde{\mathbf{q}}^{-1}K_{\lambda}^{2}-\tilde{{\bf q}}K_{\lambda}^{-2}$
are constant on $\mathcal{H}$: $Q_{\lambda}=\left(\lambda+\lambda^{-1}\right)\mathrm{id}_{\mathcal{H}}$ and $\tilde{Q}_{\lambda}=\left(\tilde{\lambda}+\tilde{\lambda}^{-1}\right)\mathrm{id}_{\mathcal{H}}$. Then $Q_{\lambda}^{\dagger}=\tilde{Q}_{\lambda}$ is equivalent to $\lambda^{*}=\tilde{\lambda}$ if we fix the “Weyl refection” $\lambda\to\lambda^{-1}$.
Alternatively, we may define
$K_{\lambda}^{\prime}=\tilde{K}_{\lambda},\qquad K_{\lambda}^{\prime-1}=\tilde{K}_{\lambda}^{-1},\qquad F_{\lambda}^{\prime}=-\tilde{E}_{\lambda},\qquad E_{\lambda}^{\prime}=-\tilde{F}_{\lambda},$
They satisfy the commutation relation of $U_{{\bf q}^{*}}(sl_{2})$
$K_{\lambda}^{\prime}F_{\lambda}^{\prime}={\bf q}^{*}F_{\lambda}^{\prime}K_{\lambda}^{\prime},\qquad K_{\lambda}^{\prime}E_{\lambda}^{\prime}={\bf q}^{*-1}E_{\lambda}^{\prime}K_{\lambda}^{\prime},\qquad\left[F_{\lambda}^{\prime},E_{\lambda}^{\prime}\right]=-\frac{\tilde{K}_{\lambda}^{2}-\tilde{K}_{\lambda}^{-2}}{{\bf q^{*}}-{\bf q}^{*-1}}.$
This shows that $\mathcal{H}$ also carries a $\star$-representation of $U_{{\bf q}}(sl_{2})\otimes U_{{\bf q}^{*}}(sl_{2})$ with
$K_{\lambda}^{\prime}=K_{\lambda}^{\dagger},\qquad E_{\lambda}^{\prime}=F_{\lambda}^{\dagger},\qquad E_{\lambda}^{\prime}=F_{\lambda}^{\dagger}.$
We may view $U_{{\bf q}}(sl_{2})\otimes U_{{\bf q}^{*}}(sl_{2})$ and $U_{{\bf q}}(sl_{2})\otimes U_{\tilde{\bf q}}(sl_{2})$ as two versions of the quantum deformation of Lorentz Lie algebra $sl_{2}\oplus sl_{2}$. $K,E,F$ and $\tilde{K},\tilde{E},\tilde{F}$ (or ${K}^{\prime},{E}^{\prime},{F}^{\prime}$) are q-deformed analogs of the self-dual and anti-self-dual generators in the Lorentz algebra. The representation $(\mathcal{H},\rho)$ of the quantum torus algebra with (3.1) and (3.2) define a family of representations of these two version of q-deformed Lorenz algebra. All the representations are infinite-dimensional and parametrized by $\lambda\in\mathbb{C}^{\times}$.
When $\mathbf{q}$ is real, $U_{{\bf q}}(sl_{2})\otimes U_{{\bf q}^{*}}(sl_{2})$ reduces to the quantum Lorentz group widely studied earlier e.g. [15, 16]. In this sense, the q-deformed Lorentz algebra considered here generalizes the quantum Lorentz group to complex $\mathbf{q}$. However, we introduce $s\in\mathbb{R}$ for parametrizing $\mathbf{q}$ by
$\mathbf{q}=\exp\left[\frac{2\pi i}{k}(1+b^{2})\right]=\exp\left[\frac{4\pi i}{k+is}\right],\qquad is=k\frac{1-b^{2}}{1+b^{2}}\ \ \text{or}\ \ b^{2}=\frac{k-is}{k+is}.$ (3.3)
The real $\mathbf{q}$ for the quantum Lorentz group corresponds to the limit $k\to 0$ and $b^{2}\to-1$ such that $s$ is finite. The representation discussed above is singular in this limit, so it is not clear how to compare the above representation with the representation of quantum Lorentz group.

### v2

@@ 2 Quantum torus algebra and the representation at level $k=2N$
Lemma 2.1.
The representation $\rho_{0}$ is irreducible in the sense that any bounded operator $\bm{\mathcal{O}}\in\mathcal{L}(\mathcal{H}_{0})$ commuting with all elements in $\mathcal{A}_{h}$ (i.e. $\bm{\mathcal{O}}\bm{a}\psi=\bm{a}\bm{\mathcal{O}}\psi$  [^2] The commutativity may be written as $\bm{\mathcal{O}}\bm{a}\subset\bm{a}\bm{\mathcal{O}}$ (the graph of $\bm{a}\bm{\mathcal{O}}$ contains the graph of $\bm{\mathcal{O}}\bm{a}$), namely, $\bm{a}\bm{\mathcal{O}}$ is an extension of $\bm{\mathcal{O}}\bm{a}$. for all $\psi\in\mathfrak{D}$ and $\bm{a}\in\mathcal{A}_{h}$) is a scalar multiple of identity operator.
Proof.
It is sufficient to show that any bounded operator on $\mathcal{H}$ commuting with $\bm{x}=\bm{x}_{1,0},\ \bm{y}=\bm{x}_{0,1},\ \widetilde{\bm{x}}=\widetilde{\bm{x}}_{1,0},\ \widetilde{\bm{y}}=\widetilde{\bm{x}}_{0,1}$ is a scalar multiple of the identity operator. The spectral decomposition of the normal operators $\bm{y},\widetilde{\bm{y}}$ gives the spectral projections $E(\Delta,m_{0}^{\prime}):L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}\to L^{2}(\Delta)$ defined by $E(\Delta,m^{\prime}_{0})f(\mu_{0},m_{0})=\chi_{\Delta}(\mu_{0})\delta_{m_{0}^{\prime}}(m_{0})f(\mu_{0},m_{0})$ for any closed interval $\Delta\subset\mathbb{R}$ and $m_{0}^{\prime}\in\mathbb{Z}/k\mathbb{Z}$. $\chi_{\Delta}$ is the characteristic function of $\Delta$. That $\bm{\mathcal{O}}$ commutes with $\bm{y},\widetilde{\bm{y}}$ implies $[\bm{\mathcal{O}},E(\Delta,m^{\prime}_{0})]=0$, so $\mathcal{O}$ leaves $L^{2}(\Delta)$ invariant. We define
$\mathcal{O}_{\Delta,m_{0}^{\prime}}(\mu_{0},m_{0}):=\left[\bm{\mathcal{O}}\left(\chi_{\Delta}\otimes\delta_{m^{\prime}_{0}}\right)\right](\mu_{0},m_{0})\in L^{2}(\Delta).$ (2.12)
$\mathcal{O}_{\Delta,m^{\prime}_{0}}(\mu_{0},m_{0})=0$ if $m_{0}\neq m^{\prime}_{0}$. For any $\Delta^{\prime}\subset\Delta$, $\bm{\mathcal{O}}E(\Delta^{\prime},m^{\prime}_{0})E(\Delta,m^{\prime}_{0})=E(\Delta^{\prime},m^{\prime}_{0})\bm{\mathcal{O}}E(\Delta,m^{\prime}_{0})$ implies $\bm{\mathcal{O}}(\chi_{\Delta^{\prime}}\otimes\delta_{m^{\prime}_{0}})=\chi_{\Delta^{\prime}}\bm{\mathcal{O}}(\chi_{\Delta}\otimes\delta_{m^{\prime}_{0}})$ and thus $\mathcal{O}_{\Delta^{\prime},m^{\prime}_{0}}(\mu_{0},m_{0})=\mathcal{O}_{\Delta,m_{0}^{\prime}}(\mu_{0},m_{0})$ for $\mu_{0}\in\Delta^{\prime}$, i.e. the value of $\mathcal{O}_{\Delta,m_{0}^{\prime}}(\mu_{0},m_{0})$ at any $\mu_{0}\in\Delta$ is independent of the choice of $\Delta$. By using any cover of $\mathbb{R}$ with closed intervals, we obtain the function $\mathcal{O}_{m_{0}^{\prime}}(\mu_{0},m_{0})$ such that $\mathcal{O}_{\Delta,m_{0}^{\prime}}(\mu_{0},m_{0})=\mathcal{O}_{m^{\prime}_{0}}(\mu_{0},m_{0})\chi_{\Delta}(\mu_{0})=\mathcal{O}_{m_{0}^{\prime}}(\mu_{0},m_{0})\chi_{\Delta}(\mu_{0})\delta_{m_{0}^{\prime}}(m_{0})$ for any $\Delta$. We define $\mathcal{O}(\mu_{0},m_{0}):=\mathcal{O}_{m_{0}}(\mu_{0},m_{0})$ and have $\mathcal{O}_{\Delta,m_{0}^{\prime}}(\mu_{0},m_{0})=\mathcal{O}(\mu_{0},m_{0})\chi_{\Delta}(\mu_{0})\delta_{m_{0}^{\prime}}(m_{0})$. For any simple function $f_{\mathfrak{p}}=\sum_{\Delta}\sum_{m_{0}}f_{\Delta,m_{0}}(\chi_{\Delta}\otimes\delta_{m_{0}})$ based on a set $\mathfrak{p}$ of disjoint $\Delta$, we obtain
$\left(\bm{\mathcal{O}}f_{\mathfrak{p}}\right)(\mu_{0},m_{0})=\sum_{\Delta\in\mathfrak{p}}\sum_{m^{\prime}_{0}\in\mathbb{Z}/k\mathbb{Z}}f_{\Delta,m^{\prime}_{0}}\mathcal{O}_{\Delta,m_{0}^{\prime}}(\mu_{0},m_{0})=\mathcal{O}(\mu_{0},m_{0})f_{\mathfrak{p}}(\mu_{0},m_{0}).$ (2.13)
For any function $f(\mu_{0},m_{0})\in\mathcal{H}$, there exists a sequence of simple functions $f_{\mathfrak{p}}$ that converges to $f$ in the sense of the Hilbert space norm. Since $\bm{\mathcal{O}}$ is a bounded operator, $\bm{\mathcal{O}}f_{\mathfrak{p}}$ converges to $\bm{\mathcal{O}}f$. Therefore,
$\left(\bm{\mathcal{O}}f\right)(\mu_{0},m_{0})=\mathcal{O}(\mu_{0},m_{0})f(\mu_{0},m_{0}),$ (2.14)
for any $f\in\mathcal{H}$. The assumption that $\bm{\mathcal{O}}\psi\in\mathfrak{D}$ for any $\psi\in\mathfrak{D}$ (in order that $\bm{a}\bm{\mathcal{O}}\psi$ is well-defined) implies that $\mathcal{O}(\mu,m)$ is an entire function in $\mu$ after analytic continuation. $|\mathcal{O}(\mu_{0},m_{0})|$ is a bounded function, since $\bm{\mathcal{O}}$ is a bounded operator.
The commutativity between $\bm{\mathcal{O}}$ and $\bm{x},\widetilde{\bm{x}}$ implies that $\mathcal{O}(\mu,m)$ is a constant: We consider $\mathcal{O}(\mu,m)$ as a tempered distribution: $\mathcal{O}\in\mathscr{S}^{\prime}(\mathbb{R})\otimes\mathbb{C}^{k}$ and for any $f\in\mathfrak{D}$,
$\mathcal{O}[f]=\sum_{m_{0}\in\mathbb{Z}/k\mathbb{Z}}\int\mathrm{d}\mu_{0}\,\mathcal{O}(\mu_{0},m_{0})f(\mu_{0},m_{0})=\sum_{n_{0}\in\mathbb{Z}/k\mathbb{Z}}\int\mathrm{d}\nu_{0}\,\mathcal{F}[\mathcal{O}](\nu_{0},n_{0})\tilde{f}(\nu_{0},n_{0}),$ (2.15)
where the Fourier transformations of the tempered distribution $\mathcal{O}$ is given by
[…]
That $\bm{x},\widetilde{\bm{x}}$ commute with $\bm{\mathcal{O}}$ implies that $e^{\frac{2\pi i}{k}\left(-ib\nu_{0}-n_{0}\right)}\mathcal{F}[\mathcal{O}](\nu_{0},n_{0})=e^{\frac{2\pi i}{k}\left(-ib^{-1}\nu_{0}+n_{0}\right)}\mathcal{F}[\mathcal{O}](\nu_{0},n_{0})=\mathcal{F}[\mathcal{O}](\nu_{0},n_{0})$, which implies $\mathcal{F}\left[\mathcal{O}\right](\nu_{0},n_{0})=\text{constant}\times\delta(\nu_{0})\delta_{\exp\left(\frac{2\pi i}{k}n_{0}\right),1}$, then $\mathcal{O}(\mu_{0},m_{0})$ is constant.
When $k=2N$, it is useful to consider the $\mathbf{q}^{2},\tilde{\mathbf{q}}^{2}$-Weyl algebra generated by $\bm{y}$, $\bm{x}^{2}\equiv\bm{u}$, $\tilde{\bm{y}}$, $\tilde{\bm{x}}^{2}\equiv\tilde{\bm{u}}$
$\bm{u}\bm{y}=\mathbf{q}^{2}\bm{y}\bm{u},\qquad\tilde{\bm{u}}\tilde{\bm{y}}=\tilde{\mathbf{q}}^{2}\tilde{\bm{y}}\tilde{\bm{u}},\qquad\bm{u}\tilde{\bm{y}}=\tilde{\bm{y}}\bm{u},\qquad\tilde{\bm{u}}{\bm{y}}={\bm{y}}\tilde{\bm{u}}.$ (2.17)
with $\mathbf{q}^{2}=e^{\frac{2\pi i}{N}\left(1+b^{2}\right)}$ and $\tilde{\mathbf{q}}^{2}=e^{\frac{2\pi i}{N}\left(1+b^{-2}\right)}$. For $k=2N$, the operator $e^{i\pi\bm{m}_{0}}$ is well-defined on $\mathcal{H}$ and commutes with the $\mathbf{q}^{2},\tilde{\mathbf{q}}^{2}$-Weyl algebra. We define the subspace $\mathcal{H}\subset\mathcal{H}_{0}$ be the eigenspace by $e^{i\pi\bm{m}_{0}}f=f$, for all $f\in\mathcal{H}$. The functions $f(\mu_{0},m_{0})$ in $\mathcal{H}$ are only supported at even $m_{0}$. If we change variables $\mu=\mu_{0}/2$ and $m=m_{0}/2$ and represent states in $\mathcal{H}$ by $\psi(\mu,m)$, we have $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$ with the inner product $\langle f,f^{\prime}\rangle=\sum_{m\in\mathbb{Z}/N\mathbb{Z}}\int\mathrm{d}\mu f(\mu,m)^{*}f^{\prime}(\mu,m)$ and
$\bm{y}\psi(\mu,m)$ $=e^{\frac{2\pi i}{N}(-ib{\mu}-{m})}\psi(\mu,m),\qquad$ $\bm{u}\psi(\mu,m)=\psi(\mu+ib,m-1),$ (2.18)
$\tilde{\bm{y}}\psi(\mu,m)$ $=e^{\frac{2\pi i}{N}(-ib^{-1}{\mu}+{m})}\psi(\mu,m),\qquad$ $\tilde{\bm{u}}\psi(\mu,m)=\psi(\mu+ib^{-1},m+1).$ (2.19)
with $\mu\in\mathbb{R}$ and $m\in\mathbb{Z}/N\mathbb{Z}$. Analogous to $\mathfrak{D}_{0}\subset\mathcal{H}_{0}$, the common dense domain $\mathfrak{D}\subset\mathcal{H}$ of $\bm{y}$,$\bm{u}$,$\tilde{\bm{y}}$,$\tilde{\bm{u}}$ and their Laurent polynomials contains $f(\mu,m)$ being entire functions in $\mu$ and satisfying
$e^{\alpha_{1}\frac{2\pi}{N}b\mu}e^{\alpha_{2}\frac{2\pi}{N}b^{-1}\mu}f(\mu+ib\beta_{1}+ib^{-1}\beta_{2},m)\in L^{2}(\mathbb{R}),\qquad\forall\,m\in\mathbb{Z}/N\mathbb{Z},\quad\alpha_{i},\beta_{i}\in\mathbb{Z}.$ (2.20)
The following operators
$\bm{u}_{\alpha\beta}=\mathbf{q}^{-\alpha\beta}\bm{u}^{\alpha}\bm{y}^{\beta}=e^{\alpha\bm{U}+\beta\bm{Y}},\qquad\tilde{\bm{u}}_{\alpha\beta}=\tilde{\mathbf{q}}^{-\alpha\beta}\tilde{\bm{u}}^{\alpha}\tilde{\bm{y}}^{\beta}=e^{\alpha\tilde{\bm{U}}+\beta\tilde{\bm{Y}}},$ (2.21)
define a representation $\rho$: $\mathcal{A}_{2\hbar}\to\mathcal{L}(\mathfrak{D})$ of the quantum torus algebra $\mathcal{A}_{2\hbar}$ with the relations:
$\bm{u}_{\alpha,\beta}\bm{u}_{\gamma,\delta}=\mathbf{q}^{\alpha\delta-\beta\gamma}\bm{u}_{\alpha+\gamma,\beta+\delta},\qquad\tilde{\bm{u}}_{\alpha,\beta}\tilde{\bm{u}}_{\gamma,\delta}=\tilde{\mathbf{q}}^{\alpha\delta-\beta\gamma}\tilde{\bm{u}}_{\alpha+\gamma,\beta+\delta}.$ (2.22)
The representation $(\mathcal{H},\rho)$ is irreducible by the same argument as in the proof of Lemma 2.1.
3 Representation of a quantum deformed Lorentz algebra
@@ 3.1 The representation
Based on the representation $(\mathcal{H},\rho)$ of the quantum torus algebra, we obtain a representation of $\mathscr{U}_{\mathbf{q}}(sl_{2})$ on $\mathcal{H}$: We define a family of operators $E_{\lambda},F_{\lambda},K_{\lambda},K_{\lambda}^{-1}$ labelled by a parameter $\lambda\in\mathbb{C}^{\times}$:
$K_{\lambda}$ $=$ $\bm{u}^{-1},\qquad K_{\lambda}^{-1}=\bm{u},\qquad F_{\lambda}=-\frac{i}{\mathbf{q}-\mathbf{q}^{-1}}q\bm{y},$ (3.1)
$E_{\lambda}$ $=$ $-\frac{i\mathbf{q}^{-1}}{\mathbf{q}-\mathbf{q}^{-1}}\left(1+\mathbf{q}\lambda^{-1}\bm{u}\right)\left(1+\mathbf{q}\lambda\bm{u}\right)q^{-1}\bm{u}^{-1}\bm{y}^{-1},$ (3.2)
@@
It is straight-forward to check that the operator algebra of $E_{\lambda},F_{\lambda},K_{\lambda},K_{\lambda}^{-1}$ represents the commutation relation of $\mathscr{U}_{\mathbf{q}}(sl_{2})$ [23]:
$KE$ $=$ $\mathbf{q}^{2}EK,\qquad KF=\mathbf{q}^{-2}FK,\qquad\left[E,F\right]=\frac{K-K^{-1}}{\mathbf{q}-\mathbf{q}^{-1}}.$ (3.3)
@@
Similarly, we define the tilded operators with $\tilde{\lambda}\in\mathbb{C}^{\times}$
$\tilde{K}_{\lambda}$ $=$ $\tilde{\bm{u}}^{-1},\qquad\tilde{K}_{\lambda}^{-1}=\tilde{\bm{u}},\qquad\tilde{F}_{\lambda}=-\frac{i}{\tilde{\mathbf{q}}-\tilde{\mathbf{q}}^{-1}}\tilde{q}^{-1}\tilde{\bm{y}},$ (3.4)
$\tilde{E}_{\lambda}$ $=$ $-\frac{i}{\tilde{\mathbf{q}}-\tilde{\mathbf{q}}^{-1}}\left(1+\tilde{\mathbf{q}}\tilde{\lambda}^{-1}\tilde{\bm{u}}\right)\left(1+\tilde{\mathbf{q}}\tilde{\lambda}\tilde{\bm{u}}\right)\tilde{q}^{-1}\tilde{\bm{u}}^{-1}\tilde{\bm{y}}^{-1}.$ (3.5)
Their operator algebra represents the commutation relation of $\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$:
$\tilde{K}\tilde{E}=\tilde{{\bf q}}^{2}\tilde{E}\tilde{K},\qquad\tilde{K}\tilde{F}=\tilde{{\bf q}}^{-2}\tilde{F}\tilde{K},\qquad\left[\tilde{E},\tilde{F}\right]=\frac{\tilde{K}-\tilde{K}^{-1}}{\tilde{{\bf q}}-\tilde{{\bf q}}^{-1}}.$ (3.6)
The tensor product $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ has the *-structure:
$E^{*}=\tilde{E},\qquad F^{*}=\tilde{F},\qquad K^{\pm 1}{}^{*}=\tilde{K}^{\pm 1},$ (3.7)
which is represented by the Hermitian conjugate on $\mathcal{H}$ when $\lambda^{*}=\tilde{\lambda}$:
$E_{\lambda}^{\dagger}=\tilde{E}_{\lambda},\qquad F_{\lambda}^{\dagger}=\tilde{F}_{\lambda},\qquad K_{\lambda}^{\pm 1\dagger}=\tilde{K}_{\lambda}^{\pm 1}.$ (3.8)
It shows that $\mathcal{H}$ carries a *-representation of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ labelled by the complex parameters $(\lambda,\tilde{\lambda})$ with $\lambda^{*}=\tilde{\lambda}$. We denote this representation by $\pi_{\lambda,\tilde{\lambda}}$. It is manifest that $\pi_{\lambda,\tilde{\lambda}}=\pi_{\lambda^{-1},\tilde{\lambda}^{-1}}$.
Lemma 3.1.
The representation $(\mathcal{H},\pi_{\lambda,\tilde{\lambda}})$ is irreducible.
Proof.
The generators of the $\mathbf{q}^{2},\widetilde{\mathbf{q}}^{2}$-Weyl algebra can be recovered by $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ generators in the following way:
$\bm{u}=K_{\lambda}^{-1},\qquad\bm{y}=i(\mathbf{q}-\mathbf{q}^{-1})q^{-1}F_{\lambda},$ (3.9)
$\tilde{\bm{u}}=\tilde{K}_{\lambda}^{-1},\qquad\tilde{\bm{y}}=i(\tilde{\mathbf{q}}-\tilde{\mathbf{q}}^{-1})\tilde{q}\tilde{F}_{\lambda}.$ (3.10)
Since any bounded operator commuting with $\bm{x},\bm{y},\widetilde{\bm{x}},\widetilde{\bm{y}}$ is a scalar multiple of identity operator, the bounded operator commuting with $\pi_{\lambda,\tilde{\lambda}}[\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})]$ also has to be a scalar multiple of identity operator. Therefore, the representation $(\mathcal{H},\pi_{\lambda,\tilde{\lambda}})$ is irreducible.
$\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ is a *-Hopf algebra, with the co-product $\Delta$, antipode $S$, and co-unit $\varepsilon$ given by
$\Delta E=E\otimes K+1\otimes E,\qquad\Delta\tilde{E}=\tilde{E}\otimes\tilde{K}+1\otimes\tilde{E},$ (3.11)
$\Delta F=F\otimes 1+K^{-1}\otimes F,\qquad\Delta\tilde{F}=\tilde{F}\otimes 1+\tilde{K}^{-1}\otimes\tilde{F}$ (3.12)
$\Delta K^{\pm 1}=K^{\pm 1}\otimes K^{\pm 1},\qquad\Delta\tilde{K}^{\pm 1}=\tilde{K}^{\pm 1}\otimes\tilde{K}^{\pm 1}$ (3.13)
$S\left(K^{\pm 1}\right)=K^{\mp 1},\qquad S\left(E\right)=-EK^{-1},\qquad S\left(F\right)=-KF,$ (3.14)
$S\left(\tilde{K}^{\pm 1}\right)=\tilde{K}^{\mp 1},\qquad S\left(\tilde{E}\right)=-\tilde{E}\tilde{K}^{-1},\qquad S\left(\tilde{F}\right)=-\tilde{K}\tilde{F}$ (3.15)
$\varepsilon\left({K}^{\pm 1}\right)=\varepsilon\left(\tilde{{K}}^{\pm 1}\right)=1,\qquad\varepsilon(E)=\varepsilon(F)=\varepsilon(\tilde{E})=\varepsilon(\tilde{F})=0.$ (3.16)
@@
$Q_{\lambda}$ $=$ $-\left({\bf q}-{\bf q}^{-1}\right)^{2}E_{\lambda}F_{\lambda}-\mathbf{q}^{-1}K_{\lambda}-{\bf q}K_{\lambda}^{-1}$ (3.17)
$\tilde{Q}_{\lambda}$ $=$ $-\left(\tilde{{\bf q}}-{\tilde{\mathbf{q}}}^{-1}\right)^{2}\tilde{E}_{\lambda}\tilde{F}_{\lambda}-\tilde{\mathbf{q}}^{-1}K_{\lambda}-\tilde{{\bf q}}K_{\lambda}^{-1}$ (3.18)
are constant on $\mathcal{H}$: $Q_{\lambda}=\left(\lambda+\lambda^{-1}\right)\mathrm{id}_{\mathcal{H}}$ and $\tilde{Q}_{\lambda}=\left(\tilde{\lambda}+\tilde{\lambda}^{-1}\right)\mathrm{id}_{\mathcal{H}}$. Then $Q_{\lambda}^{\dagger}=\tilde{Q}_{\lambda}$ when $\lambda^{*}=\tilde{\lambda}$.
[…]
$[E,F]=2H,\qquad[H,E]=E,\qquad[H,F]=-F,$ (3.19)
$[\tilde{E},\tilde{F}]=2\tilde{H},\qquad[\tilde{H},\tilde{E}]=\tilde{E},\qquad[\tilde{H},\tilde{F}]=-\tilde{F},$ (3.20)
and
$E^{*}=\tilde{E},\qquad F^{*}=\tilde{F},\qquad H^{*}=-\tilde{H}.$ (3.21)
The relations (3.3) and (3.6) of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ are deformations of (3.19) and (3.20) by identifying $K=\mathbf{q}^{2H}$ and $\tilde{K}=\tilde{\mathbf{q}}^{2\tilde{H}}$. The *-structure of $\mathscr{U}_{\mathbf{q}}(sl_{2})\otimes\mathscr{U}_{\widetilde{\mathbf{q}}}(sl_{2})$ in (3.7) follows from (3.21) due to $\mathbf{q}^{*}=\tilde{\mathbf{q}}^{-1}$.